The length of a rectangle is five times its width. If the perimeter of the rectangle is 60 m , find its area.
Found 2 solutions by jedidiah, rajagopalan:
Answer by jedidiah(6)

width: x
length: 5x
12x = 60
x=5
5*25= 125 units^2

Answer by rajagopalan(174)

Let the Width = x
Then Length = 5x
Perimeter=2(x+5x)=12x
12x=60
x=60/12=5
so Length =5*5=25
Breadth=x=5
Area = L*B
=25*5
=125 Sq m
Area=125 Sq M